B + C = 24 yards. B is 6 yards shorter than C. How do we set up this problem to solve?
Answer by nerdybill(7384)

B + C = 24 yards. B is 6 yards shorter than C. How do we set up this problem to solve?
.
Because "B is 6 yards shorter than C" we have:
B = C-6
.
substitute above into the given equation:
B + C = 24
to get:
C-6 + C = 24
2C-6 = 24
2C = 30
C = 15
.
To find B, use:
B = C-6
B = 15-6
B = 9